I have one more equation that I am having problems with. Once again, the answer I came up with does not seem to work when I am trying to check it. The problem is as follows:
8z+2=2(z-5)-z
I come up with an answer of 1.7, but I cannot get the check to work out. Is my answer correct?

Answer by dolly(163)

8z+2=2(z-5)-z
Adding z to both the sides,
8z + 2 + z = 2z - 10
adding - 2z to both the sides,
8z + 2 + z - 2z = - 10
7z + 2 = - 10
adding -2 to both the sides,
7z = - 10 -2
7z = - 12
z = - 12/7